OpenStudy (anonymous):

Perform the indicated operation. (a + 3)^2/a-3 / 5/(a^2-9)

OpenStudy (stamp):

\[\frac{(a+3)^2}{a-3}\div\frac{5}{a^2-9}=\]\[\frac{(a+3)^2}{a-3}*\frac{a^2-9}{5}\]\[=\frac{(a+3)(a+3)}{(a-3)}\frac{(a+3)(a-3)}{5}\]\[=\frac{(a+3)^3}{5}\]